# DECIPHERING THE DEFINITION OF ADVERSARIAL ROBUSTNESS FOR POST-HOC OOD DETECTORS

### v2

@@ 4 EXPERIMENTS
Detector Attacks CIFAR-10 CIFAR-100 ImageNet-200 ImageNet-1K
ResNet-18 ResNet-50 Swin-T
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
SCALE PGD 99.67 34.53 99.97 16.18 95.49 35.14 100.00 0.20 95.49 35.14
FGSM 85.74 77.50 49.69 85.64 79.75 76.88 89.75 66.28 79.75 76.88
DF 67.07 81.73 69.22 68.69 79.75 76.88 87.82 57.77 79.75 76.88
mPGD 88.50 70.69 85.58 59.67 93.24 42.17 100.00 6.90 93.24 42.17
NNGUIDE PGD 99.39 30.29 98.85 17.07 96.44 33.60 100.00 0.12 96.44 33.60
FGSM 93.10 53.01 68.14 77.62 83.21 75.13 85.27 73.53 83.21 75.13
DF 92.08 63.25 85.36 64.14 83.21 75.13 82.19 62.81 83.21 75.13
mPGD 92.94 58.90 90.98 57.26 94.30 42.87 99.99 9.52 94.30 42.87
GEN PGD 99.51 41.75 99.90 26.03 89.17 40.03 100.00 0.21 89.17 40.03
FGSM 70.14 81.29 45.66 87.10 72.06 79.00 83.63 73.28 72.06 79.00
DF 44.32 85.98 71.38 65.82 72.06 79.00 80.01 62.96 72.06 79.00
mPGD 83.65 74.35 75.59 66.56 88.10 47.17 99.96 12.40 88.10 47.17
ASH PGD 99.67 31.23 99.96 24.49 97.06 32.97 100.00 0.19 97.06 32.97
FGSM 86.94 70.60 42.86 88.14 83.81 74.57 85.61 69.99 83.81 74.57
DF 77.10 74.62 74.98 65.44 83.81 74.57 83.29 60.88 83.81 74.57
mPGD 90.75 64.35 78.87 66.61 95.20 40.17 100.00 8.49 95.20 40.17
DICE PGD 96.65 36.09 99.93 23.45 95.03 34.24 100.00 0.11 95.03 34.24
FGSM 75.86 72.65 46.27 86.51 75.62 78.73 86.99 71.20 75.62 78.73
DF 68.84 73.44 76.73 65.46 75.62 78.73 84.72 62.92 75.62 78.73
mPGD 89.46 66.25 80.48 66.68 91.60 42.99 99.99 8.63 91.60 42.99
KNN PGD 64.91 69.18 90.06 43.07 85.23 55.53 78.63 55.74 85.23 55.53
FGSM 61.23 82.08 47.81 84.69 76.88 73.23 75.89 68.43 76.88 73.23
DF 38.78 85.84 78.54 63.06 76.88 73.23 86.65 58.09 76.88 73.23
mPGD 76.02 75.02 78.93 64.03 88.63 54.44 87.09 48.40 88.63 54.44
VIM PGD 92.45 56.83 98.16 42.79 89.17 46.44 100.00 5.16 89.17 46.44
FGSM 54.89 84.43 54.72 74.75 71.60 69.44 75.28 71.31 71.60 69.44
DF 43.92 84.92 78.57 61.75 71.60 69.44 82.95 59.88 71.60 69.44
mPGD 80.70 74.35 80.56 60.01 90.16 50.71 99.86 24.51 90.16 50.71
KLM PGD 91.43 60.91 91.94 45.75 90.87 54.58 95.49 40.77 90.87 54.58
FGSM 96.90 66.02 72.31 80.83 80.71 74.44 80.54 71.56 80.71 74.44
DF 80.84 71.64 91.38 59.58 80.71 74.44 85.48 59.30 80.71 74.44
mPGD 97.52 62.26 89.69 61.30 91.96 55.14 94.81 41.19 91.96 55.14
MLS PGD 99.58 39.65 99.96 24.43 94.43 35.64 100.00 0.12 94.43 35.64
FGSM 75.61 80.97 43.04 87.63 74.47 79.06 85.30 74.16 74.47 79.06
DF 51.12 84.89 74.91 65.41 74.47 79.06 81.44 63.37 74.47 79.06
mPGD 85.11 73.78 78.81 66.41 90.90 44.19 99.97 10.65 90.90 44.19
REACT PGD 98.84 45.19 99.89 25.13 94.49 35.90 100.00 4.15 94.49 35.90
FGSM 79.55 79.84 42.83 88.27 76.12 78.15 80.31 74.14 76.12 78.15
DF 54.80 84.16 74.99 65.49 76.12 78.15 79.97 62.88 76.12 78.15
mPGD 85.32 73.19 78.70 66.52 91.26 44.59 99.71 20.13 91.26 44.59
GRAM PGD 99.82 22.50 99.94 17.12 98.75 25.92 100.00 0.07 98.75 25.92
FGSM 94.77 56.54 79.42 79.23 91.60 69.07 96.06 59.67 91.60 69.07
DF 88.14 60.87 90.78 58.04 91.60 69.07 93.66 55.04 91.60 69.07
mPGD 93.21 55.56 91.49 58.72 97.83 34.92 100.00 4.36 97.83 34.92
RMDS PGD 49.03 82.70 66.28 77.08 53.69 76.40 37.47 95.23 53.69 76.40
FGSM 68.03 80.66 76.00 79.46 67.10 76.26 71.65 73.33 67.10 76.26
DF 43.26 85.75 64.52 80.96 67.10 76.26 73.74 65.11 67.10 76.26
mPGD 77.05 75.84 84.46 74.74 79.37 63.99 90.78 50.99 79.37 63.99
EBO PGD 99.58 39.61 99.96 24.49 94.47 35.41 100.00 0.12 94.47 35.41
FGSM 75.62 81.04 42.86 88.14 74.58 79.17 85.35 74.42 74.58 79.17
DF 51.23 84.71 74.99 65.44 74.58 79.17 81.63 63.79 74.58 79.17
mPGD 85.11 73.80 78.87 66.61 91.00 44.10 99.97 10.64 91.00 44.10
MDS PGD 47.81 84.48 50.82 84.18 57.34 76.11 0.05 99.95 57.34 76.11
FGSM 64.24 79.22 86.06 51.31 91.41 49.61 90.61 52.93 91.41 49.61
DF 54.04 79.60 89.76 52.53 91.41 49.61 93.19 49.18 91.41 49.61
mPGD 78.25 69.86 91.87 50.63 83.81 65.37 51.74 89.98 83.81 65.37
ODIN PGD 99.73 33.25 99.97 17.58 97.20 31.19 100.00 0.91 97.20 31.19
FGSM 73.56 83.16 39.74 90.19 77.39 76.38 85.44 72.56 77.39 76.38
DF 60.16 83.99 72.94 68.00 77.39 76.38 82.00 65.46 77.39 76.38
mPGD 86.49 74.11 85.66 62.09 91.72 45.12 99.90 15.89 91.72 45.12
MSP PGD 99.34 43.62 100.00 27.04 87.66 42.60 100.00 4.08 87.66 42.60
FGSM 67.16 81.14 47.87 85.08 72.12 77.77 80.39 70.45 72.12 77.77
DF 39.78 86.85 70.63 65.92 72.12 77.77 72.65 60.76 72.12 77.77
mPGD 82.76 74.68 75.03 66.23 87.48 49.07 100.00 22.21 87.48 49.07
@@ 4.1 DISCUSSION
All methods do not show sufficient results, when evaluating 16 post-hoc methods as unveiled in the Table 3. Only two of them - Mahalanobis distance-based methods - [64, 65] show partly detection capabilities against the FGSM, PGD, and mPGD attacks on ResNet-50 with ImageNet-1K. The Mahalanobis distance’s robustness has been thoroughly studied [66, 67, 68, 69], and its adversarial robustness is attributed to its covariance structure consideration [67]. This highlights the conflict between AML and OOD detection, where detectors often excel in either adversarial or natural distributions, but not both. High detection rates on OOD samples are the fundament for further defense mechanisms.
Level of Adversarial Robustness - From Detectors towards Defenses A defense is more sophisticated than detectors to mitigate the attacker’s efforts to fool a classifier. A step towards adversarial defense could be to improve adversarial robustness in OOD detectors. We suggest a possible roadmap to evaluate detectors and lift them toward an adversarial defense:
  1. 1.
@@
  2. 2.
    Use different models and other datasets than the simple ResNet-18 trained on CIFAR-10. We suggest using ImageNet-1K because its complexity in terms of resolution and objects is more closely to real-world scenarios.
  3. 3.
@@
  4. 4.
    Let your method fail against sophisticated attacks, such as adaptive attacks [72, 50] or design OOD adversarial examples that convert OD to ID samples [3]. Adversarial robustness is an iterative process, where defenses are proposed, evaluated, and then improved upon in response to new attacks or discovered vulnerabilities.
@@ 5 CONCLUSION
In this study, we assess the performance of the 16 post-hoc OOD detectors in their ability to detect various evasive attacks. We conducted prominent white-box adversarial attacks, such as PGD and DeepFool, on the CIFAR-10 and ImageNet-1K datasets. Our discovery indicates that current post-hoc methods are not ready for real-world applications as long as they are vulnerable against well-known threat — adversarial examples. We hope that our experiments give a baseline for further research by improving post-hoc methods towards robustness and will find a place as a standardized benchmark, such as OpenOOD.
Future Work We propose to extend the experiments towards transferability because adversarial examples transfer effectively across different datasets [73] and models [74]. Finally, we would suggest using black-box attacks for a realistic open-world scenario.
Appendix A OOD Definition per Method
This section extends the Section 3. In Table 4, we compare the chosen ID and OOD datasets of each OOD detector. In this comparison, we pick those experiments with largest dimensioned available datasets. Furthermore, we extend this comparison by appending the adversarial robust OOD detectors from the related work in Section 2.3. It can be observed that the more recent post-hoc detectors tend to have ImageNet-1K as standard ID datasets. In contrast, the adversarial robust OOD detectors are benchmarked more likely on the smaller and less complex CIFAR-10 dataset.
Following attacks, models, and datasets are used to heuristically evaluate OOD samples (more details in Table 4):
Attacks: FGSM [27], PGD [1]
Model Architectures:
  • ResNet-18/ResNet-50 [62], ResNetv2-101 [75], ResNet-50-D [76], TResNet-M [77], WideResNet [78]
  • BiT [79], VIT-B-16 [80], ViT [81], DeiT [81], Swin-T [63]
  • DenseNet-121 & DenseNet-101 [82]
  • MobileNet [83], MobileNetV2 [84]
  • RegNet & RegNetX4.0 [85]
  • RepVGG [86]
  • Mixer-B-16 [87]
  • CLIP [88]
Datasets: CIFAR-10/CIFAR-100 [18], BDD-Anomaly [55], DeepDrive [89] ImageNet-1K [19], ImageNet-21K [90], ImageNet-O [7], iNaturalist [13], ISUN [91], GSTRB [92], LSUN [93], NINCO [56], OpenImage-O [9], Pascal-VOC [94], Places [95], Textures [11], TinyImageNet [96], Species [97], StreetHazards [55], SSB-hard [98], SVHN [99], SUN [100].
Table 4: Overview of ID and OOD definition from several OOD detectors. The detectors are divided into the following categories [5]: Classifiction-based, Density-based, Distance-based. We also mark: Supervised and Adversarial Robust.
Methods ID OOD Model Architectures
PostHoc Methods
KNN [16] ImageNet-1K iNaturalist, Places, SUN, Textures ResNet-50
EBO [106] CIFAR-10 ISUN, Places, Texture, SVHN, LSUN WideResNet
ODIN [107] CIFAR-10 LSUN, SVHN, TinyImageNet DenseNet, ResNet
MSP [8] CIFAR-10 SUN (Gaussian) WideResNet 40-4
OOD Detectors for Adversarial Robustness
ALOE [21] CIFAR-10, CIFAR-100, GSTRB PGD attack DenseNet
OSAD [51] CIFAR-10, SVHN, TinyImageNet FGSM, PGD attack ResNet-18
ADT [42] CIFAR-10, CIFAR-100 FGSM, PGD attack ViT
ATOM [52] CIFAR-10, CIFAR-100, SVHN PGD attack WideResNet
SAFE [38] PASCAL-VOC, DeepDrive FGSM attack RegNetX4.0, ResNet-50

### v3

@@ 4.2 Grad-CAM Similarity
Detector Attacks CIFAR-10 CIFAR-100 ImageNet-200 ImageNet-1K ResNet-18 ResNet-50 Swin-T FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC SCALE PGD 99.67 34.53 99.97 16.18 95.49 35.14 100.00 0.20 95.49 35.14 FGSM 85.74 77.50 49.69 85.64 79.75 76.88 89.75 66.28 79.75 76.88 DF 67.07 81.73 69.22 68.69 79.75 76.88 87.82 57.77 79.75 76.88 mPGD 88.50 70.69 85.58 59.67 93.24 42.17 100.00 6.90 93.24 42.17 NNGUIDE PGD 99.39 30.29 98.85 17.07 96.44 33.60 100.00 0.12 96.44 33.60 FGSM 93.10 53.01 68.14 77.62 83.21 75.13 85.27 73.53 83.21 75.13 DF 92.08 63.25 85.36 64.14 83.21 75.13 82.19 62.81 83.21 75.13 mPGD 92.94 58.90 90.98 57.26 94.30 42.87 99.99 9.52 94.30 42.87 GEN PGD 99.51 41.75 99.90 26.03 89.17 40.03 100.00 0.21 89.17 40.03 FGSM 70.14 81.29 45.66 87.10 72.06 79.00 83.63 73.28 72.06 79.00 DF 44.32 85.98 71.38 65.82 72.06 79.00 80.01 62.96 72.06 79.00 mPGD 83.65 74.35 75.59 66.56 88.10 47.17 99.96 12.40 88.10 47.17 ASH PGD 99.67 31.23 99.96 24.49 97.06 32.97 100.00 0.19 97.06 32.97 FGSM 86.94 70.60 42.86 88.14 83.81 74.57 85.61 69.99 83.81 74.57 DF 77.10 74.62 74.98 65.44 83.81 74.57 83.29 60.88 83.81 74.57 mPGD 90.75 64.35 78.87 66.61 95.20 40.17 100.00 8.49 95.20 40.17 DICE PGD 96.65 36.09 99.93 23.45 95.03 34.24 100.00 0.11 95.03 34.24 FGSM 75.86 72.65 46.27 86.51 75.62 78.73 86.99 71.20 75.62 78.73 DF 68.84 73.44 76.73 65.46 75.62 78.73 84.72 62.92 75.62 78.73 mPGD 89.46 66.25 80.48 66.68 91.60 42.99 99.99 8.63 91.60 42.99 KNN PGD 64.91 69.18 90.06 43.07 85.23 55.53 78.63 55.74 85.23 55.53 FGSM 61.23 82.08 47.81 84.69 76.88 73.23 75.89 68.43 76.88 73.23 DF 38.78 85.84 78.54 63.06 76.88 73.23 86.65 58.09 76.88 73.23 mPGD 76.02 75.02 78.93 64.03 88.63 54.44 87.09 48.40 88.63 54.44 VIM PGD 92.45 56.83 98.16 42.79 89.17 46.44 100.00 5.16 89.17 46.44 FGSM 54.89 84.43 54.72 74.75 71.60 69.44 75.28 71.31 71.60 69.44 DF 43.92 84.92 78.57 61.75 71.60 69.44 82.95 59.88 71.60 69.44 mPGD 80.70 74.35 80.56 60.01 90.16 50.71 99.86 24.51 90.16 50.71 KLM PGD 91.43 60.91 91.94 45.75 90.87 54.58 95.49 40.77 90.87 54.58 FGSM 96.90 66.02 72.31 80.83 80.71 74.44 80.54 71.56 80.71 74.44 DF 80.84 71.64 91.38 59.58 80.71 74.44 85.48 59.30 80.71 74.44 mPGD 97.52 62.26 89.69 61.30 91.96 55.14 94.81 41.19 91.96 55.14 MLS PGD 99.58 39.65 99.96 24.43 94.43 35.64 100.00 0.12 94.43 35.64 FGSM 75.61 80.97 43.04 87.63 74.47 79.06 85.30 74.16 74.47 79.06 DF 51.12 84.89 74.91 65.41 74.47 79.06 81.44 63.37 74.47 79.06 mPGD 85.11 73.78 78.81 66.41 90.90 44.19 99.97 10.65 90.90 44.19 REACT PGD 98.84 45.19 99.89 25.13 94.49 35.90 100.00 4.15 94.49 35.90 FGSM 79.55 79.84 42.83 88.27 76.12 78.15 80.31 74.14 76.12 78.15 DF 54.80 84.16 74.99 65.49 76.12 78.15 79.97 62.88 76.12 78.15 mPGD 85.32 73.19 78.70 66.52 91.26 44.59 99.71 20.13 91.26 44.59 GRAM PGD 99.82 22.50 99.94 17.12 98.75 25.92 100.00 0.07 98.75 25.92 FGSM 94.77 56.54 79.42 79.23 91.60 69.07 96.06 59.67 91.60 69.07 DF 88.14 60.87 90.78 58.04 91.60 69.07 93.66 55.04 91.60 69.07 mPGD 93.21 55.56 91.49 58.72 97.83 34.92 100.00 4.36 97.83 34.92 RMDS PGD 49.03 82.70 66.28 77.08 53.69 76.40 37.47 95.23 53.69 76.40 FGSM 68.03 80.66 76.00 79.46 67.10 76.26 71.65 73.33 67.10 76.26 DF 43.26 85.75 64.52 80.96 67.10 76.26 73.74 65.11 67.10 76.26 mPGD 77.05 75.84 84.46 74.74 79.37 63.99 90.78 50.99 79.37 63.99 EBO PGD 99.58 39.61 99.96 24.49 94.47 35.41 100.00 0.12 94.47 35.41 FGSM 75.62 81.04 42.86 88.14 74.58 79.17 85.35 74.42 74.58 79.17 DF 51.23 84.71 74.99 65.44 74.58 79.17 81.63 63.79 74.58 79.17 mPGD 85.11 73.80 78.87 66.61 91.00 44.10 99.97 10.64 91.00 44.10 MDS PGD 47.81 84.48 50.82 84.18 57.34 76.11 0.05 99.95 57.34 76.11 FGSM 64.24 79.22 86.06 51.31 91.41 49.61 90.61 52.93 91.41 49.61 DF 54.04 79.60 89.76 52.53 91.41 49.61 93.19 49.18 91.41 49.61 mPGD 78.25 69.86 91.87 50.63 83.81 65.37 51.74 89.98 83.81 65.37 ODIN PGD 99.73 33.25 99.97 17.58 97.20 31.19 100.00 0.91 97.20 31.19 FGSM 73.56 83.16 39.74 90.19 77.39 76.38 85.44 72.56 77.39 76.38 DF 60.16 83.99 72.94 68.00 77.39 76.38 82.00 65.46 77.39 76.38 mPGD 86.49 74.11 85.66 62.09 91.72 45.12 99.90 15.89 91.72 45.12 MSP PGD 99.34 43.62 100.00 27.04 87.66 42.60 100.00 4.08 87.66 42.60 FGSM 67.16 81.14 47.87 85.08 72.12 77.77 80.39 70.45 72.12 77.77 DF 39.78 86.85 70.63 65.92 72.12 77.77 72.65 60.76 72.12 77.77 mPGD 82.76 74.68 75.03 66.23 87.48 49.07 100.00 22.21 87.48 49.07
@@ 4.3 DISCUSSION
All 16 post-hoc methods evaluated in our study as shown in Table 4 demonstrate inadequate performance. Notably, only two methods based on Mahalanobis distance [11, 74] exhibit partial detection capabilities against FGSM, PGD, and mPGD attacks on ResNet-50 with ImageNet-1K. The robustness of Mahalanobis distance has been extensively explored [87, 88, 89, 90], attributing its resilience to adversarial attacks to its consideration of covariance structures [88].
Interestingly, several OOD benchmarks [91, 18, 5, 6] and a recent challenge [92] focus solely on natural distribution shifts, neglecting adversarial examples. This underscores a fundamental conflict between Adversarial Machine Learning (AML) and Out-of-Distribution (OOD) detection, where detectors excel either in adversarial or natural distributions, but rarely both. Current OOD detectors struggle to converge on both fronts.
Adversarial examples, as counterparts of In-Distribution (ID) data, significantly alter the attention of pre-trained classifiers, as shown in Figure 1 and discussed in Equation 3. These predominantly untargeted attacks aim to minimally change class predictions, resulting in dissimilarity in attention maps (see Grad-CAM Section 4.2). The DF attack exhibits minimal attention shifts across numerous samples, yet many attacked samples still demonstrate notable deviations in attention maps. Similarly, the mPGD attack induces perceptible perturbations, with slightly improved detection results compared to PGD, but overall performance remains insufficient.
Post-hoc detectors are trained on distributions similar to the underlying pre-trained model, yet studies (e.g., [93, 94]) indicate significant discrepancies between the training distribution and adversarial distributions. Achieving high detection rates on adversarial OOD samples is crucial for developing robust defense mechanisms, a goal that remains elusive in current research.
Levels of Adversarial Robustness - From Detector towards Defense
A defense is more sophisticated than detectors to mitigate attackers’ efforts to fool a classifier. A step towards adversarial defense could be to improve adversarial robustness in OOD detectors. We suggest a possible roadmap to evaluate detectors and lift them toward an adversarial defense on different levels:
  1. Level 1:
@@
  2. Level 2:
    Use different models and datasets than the simple ResNet-18 trained on CIFAR-10. We suggest using ImageNet-1K because its resolution and objects are more closely related to real-world scenarios.
  3. Level 3:
@@
  4. Level 4:
    Let your method fail against sophisticated attacks, such as adaptive attacks [97, 54] or design OOD adversarial examples that convert OD to ID samples [3]. Adversarial robustness is an iterative process, where defenses are proposed, evaluated, and improved upon in response to new attacks or discovered vulnerabilities.
It must be noted that not all levels are required to be evaluated. This roadmap should indicate which level of OOD detection is robust. Further research would be encouraged to do at least an adversarial robustness benchmark (level 1), even though the OOD detector is not focused on adversarial robustness.
@@ 5 CONCLUSION
In this study, we assess the performance of the 16 post-hoc OOD detectors in their ability to detect various evasive attacks. We conducted prominent white-box adversarial attacks, such as PGD and DF, on the CIFAR-10 and ImageNet-1K datasets. Our discovery indicates that current post-hoc methods are not ready for real-world applications as long as they are vulnerable against the well-known threat — adversarial examples. We hope our experiments give a baseline for further research by improving post-hoc methods towards robustness and will find a place as a standardized benchmark, such as OpenOOD.
Future Work
We propose to extend the experiments towards transferability because adversarial examples transfer effectively across different datasets [98] and models [99]. Then, we would suggest using black-box attacks for a realistic open-world scenario. Throughout this work, we assumed a perfect pre-trained classifier on a clean curated dataset. An imperfect pre-trained classifier would be the subject of future research [100].